# The Relational Bottleneck as an Inductive Bias for Efficient Abstraction

### v2

@@ Abstract
Box 2: Neuro-symbolic modeling approaches
More recently, hybrid systems have been developed that combine deep learning with symbolic programs. In this approach, deep learning components are typically employed to translate raw perceptual inputs, such as images or natural language, into symbolic representations, which can then be processed by traditional symbolic algorithms [64, 65, 66, 67, 68]. This approach is complemented by recent neuro-symbolic approaches to probabilistic program induction, in which symbolic primitives are pre-specified (following earlier symbolic-connectionist modeling efforts), and then deep learning is used to assemble these primitives into programs [17].
@@
To address this, a recent study investigated the development of counting in deep neural network architectures [40]. These included the ESBN, the Transformer, and long short-term memory (LSTM) [81] (a type of recurrent neural network). Each architecture displayed a distinct developmental timecourse (Figure 3). The Transformer displayed a roughly linear timecourse, taking approximately the same amount of time to master each number. The LSTM displayed an exponentially increasing timecourse, taking more time to learn each new number. Only the ESBN displayed a human-like inductive transition, gradually learning to count each number from one to four, and then rapidly acquiring the ability to count higher after learning to count to five. This was due to the ability of the ESBN to learn a procedure over the representations in its control pathway that was abstracted away from the specific numbers in the count sequence (represented in the model’s perceptual pathway), allowing it to rapidly and systematically generalize between numbers. Specifically, the ESBN learned a procedure in which it stopped counting once the count sequence matched the desired number of objects (once the target value of N had been reached in the give-N task). The ESBN’s use of symbol-like representations (in its control pathway, representing ‘item at which to stop’) allowed this procedure to be abstracted away from the particular target value (represented in the perceptual pathway, and bound to the symbol-like representation in external memory), facilitating rapid generalization to higher values. This case study illustrates how the relational bottleneck can facilitate a human-like developmental trajectory for learning abstract concepts.
@@
The relational bottleneck principle may also help to explain the limited capacity of some cognitive processes (e.g., working memory) [82]. Recent work has demonstrated that human-like capacity limits naturally emerge in an architecture that implements the relational bottleneck [83]. In that architecture, two separate representational pools (each representing distinct feature spaces, e.g., color and location) interact via a dynamic variable-binding mechanism (in that case, implemented using rapid Hebbian learning [84]) in a manner that is conceptually similar to the ESBN. This mechanism enables the model to flexibly construct compositional representations (e.g., representing a visual scene by binding together spatial locations and visual features). However, this flexibility comes at the cost of relying on compositional representations that, by definition, are shared across many different, potentially competing processes (an instance of the general relationship between shared representations and cognitive capacity [85]), a problem often referred to in psychology as the ‘binding problem’ [86]. For instance, the representation of a blue object in the upper left corner (formed by binding together representations of ‘blue’ and ‘upper left’) will overlap with the representation of a blue object in the lower right corner (formed by binding together representations of ‘blue’ and ‘lower right’), leading to interference between these two representations (note that this interference is also further exacerbated by the use of maximally low-dimensional codes to represent features). The model quantitatively captures capacity limits observed in three distinct cognitive domains: working memory [82], subitizing (the ability to rapidly identify the number of items in a display) [87], and absolute judgment (the ability to correctly label specific feature values such as pitch or loudness) [88].
@@
However, this segregation raises the question of how representations in these distinct neocortical systems are flexibly bound together. Though many proposals have been made for how the brain might solve this variable-binding problem (see Box 2), one intriguing possibility involves use of the episodic memory (EM) system [44]. A common view holds that EM is supported by rapid synaptic plasticity in the hippocampus, which complements slower statistical learning in the neocortex [45, 97]. According to this view, episodes are encoded in the hippocampus by the rapid binding of features that co-occur within an episode, while the features themselves are represented in neocortical systems. This same mechanism could in principle support an architecture similar to the ESBN, by enabling rapid binding of abstract and perceptual neocortical representations. This is in fact very similar to models of cognitive map learning, in which conjunctive representations are rapidly formed in the hippocampus [98]. These hippocampal conjunctive codes bind structural and sensory information, which are thought to be encoded in the medial and lateral entorhinal cortices, respectively, and are commonly understood as extensions of the parietal and temporal neocortical systems referenced above. More generally, the involvement of EM in relational reasoning would be consistent with a growing body of recent findings suggesting the potential involvement of EM in tasks traditionally associated with working memory [99, 100, 101].
@@
There are a number of important avenues for further developing the relational bottleneck framework (see Outstanding Questions). Further work is needed to integrate the relational bottleneck with a broader range of cognitive processes relevant to abstraction, including attentional processes [121] and semantic cognition [122]. Additionally, much work has suggested that human reasoning is not purely relational, but instead depends on a mixture of concrete and abstract influences [123, 124, 125, 126]. This suggests the potential value of a more graded formulation that controls the amount of non-relational information allowed to pass through the bottleneck. Finally, the human capacity for abstraction surely depends not only on architectural biases such as those that we have discussed here, but also on the rich educational and cultural fabric that allows us to build on the abstractions developed by others [127]. In future work, it will be important to explore the interaction between education, culture and relational inductive biases.
@@
  • Human reasoners often display so-called ‘content effects’, in which abstract reasoning processes are influenced by the specific content under consideration (and therefore are not purely abstract or relational). Can a more graded version of the relational bottleneck capture these effects, while preserving a capacity for relational abstraction?
@@
  • The relational bottleneck principle affords an especially important role to patterns of similarity, mirroring the central role played by similarity in some cognitive models of analogical reasoning. What is the relationship between these modeling approaches? Does the relational bottleneck provide a useful inductive bias toward learning to implement processes such as analogical mapping?
  • How can other cognitive processes (perception, attention, memory, etc.) be integrated with the relational bottleneck?
@@
  • How do architectural biases toward relational processing interact with cultural sources of abstraction (e.g., formal education)?
@@
In the context of deep learning, an approach that combines neural networks with separate external stores of information, typically with learnable mechanisms for writing to and reading from these stores, and in which retrieval is usually similarity-based (i.e, ‘content-addressable’). Often used to implement a form of episodic memory [131].

### v5

@@ Abstract
To address this, a recent study investigated the development of counting in deep neural network architectures [56]. These included the ESBN, the Transformer, and long short-term memory (LSTM) [80] (a type of recurrent neural network). Each architecture displayed a distinct developmental timecourse (Figure 3(c)). The Transformer displayed a roughly linear timecourse, taking approximately the same amount of time to master each number. The LSTM displayed an exponentially increasing timecourse, taking more time to learn each new number. Only the ESBN displayed a human-like inductive transition, gradually learning to count each number from one to four, and then rapidly acquiring the ability to count higher after learning to count to five. This was due to the ability of the ESBN to learn a procedure over the representations in its control pathway that was abstracted away from the specific numbers in the count sequence (represented in the model’s perceptual pathway), allowing it to rapidly and systematically generalize between numbers. Specifically, the ESBN learned a procedure in which it stopped counting once the count sequence matched the desired number of objects (once the target value of N had been reached in the give-N task). The ESBN’s use of symbol-like representations (in its control pathway, representing ‘item at which to stop’) allowed this procedure to be abstracted away from the particular target value (represented in the perceptual pathway, and bound to the symbol-like representation in external memory), facilitating rapid generalization to higher values. This case study illustrates how the relational bottleneck can facilitate a human-like developmental trajectory for learning abstract concepts.
@@
The relational bottleneck principle may also help to explain the limited capacity of some cognitive processes (e.g., working memory) [81]. Recent work has proposed a normative explanation of capacity-limited processes, according to which these capacity limits arise from the use of compositional representations, implemented in an architecture that employs a relational bottleneck [82]. In that architecture, two separate representational pools (each representing distinct feature spaces, e.g., color and location) interact via a dynamic variable-binding mechanism (in that case, implemented using rapid Hebbian learning [83]) in a manner that is conceptually similar to the ESBN. This mechanism enables the model to flexibly construct compositional representations (e.g., representing a visual scene by binding together spatial locations and visual features). However, this flexibility comes at the cost of relying on compositional representations that, by definition, are shared across many different, potentially competing processes. For instance, the representation of a blue object in the upper left corner (formed by binding together representations of ‘blue’ and ‘upper left’) will overlap with the representation of a blue object in the lower right corner (formed by binding together representations of ‘blue’ and ‘lower right’), leading to interference between these two representations (note that this interference is also further exacerbated by the use of maximally low-dimensional codes to represent features). This can be viewed as an instance of the more general tradeoff between processing capacity and the use of shared representations [84], and also relates to classic considerations regarding the ‘binding problem’ in psychology [85], though it has not been previously appreciated that this tradeoff can explain the severely capacity-limited nature of processes such as working memory (and others, including subitizing [86], and absolute judgment [87]). This work illustrates how one of the most notable strengths of human cognition – compositionality – may explain one of its most notable weaknesses – capacity-limited processing – and how both can be implemented in neural networks via the relational bottleneck.
@@
However, this segregation raises the question of how representations in these distinct neocortical systems are flexibly bound together. Though many proposals have been made for how the brain might solve this variable-binding problem (see Box 1), one intriguing possibility involves use of the episodic memory (EM) system [60]. A common view holds that EM is supported by rapid synaptic plasticity in the hippocampus, which complements slower statistical learning in the neocortex [61, 96]. According to this view, episodes are encoded in the hippocampus by the rapid binding of features that co-occur within an episode, while the features themselves are represented in neocortical systems. This same mechanism could in principle support an architecture similar to the ESBN, by enabling rapid binding of abstract and perceptual neocortical representations. This is in fact very similar to models of cognitive map learning, in which conjunctive representations are rapidly formed in the hippocampus [97]. These hippocampal conjunctive codes bind structural and sensory information, which are thought to be encoded in the medial and lateral entorhinal cortices, respectively, and are commonly understood as extensions of the parietal and temporal neocortical systems referenced above. More generally, the involvement of EM in relational reasoning would be consistent with a growing body of recent findings suggesting the potential involvement of EM in tasks traditionally associated with working memory [98, 99, 100].
@@
There are a number of important avenues for further developing the relational bottleneck framework. One major question concerns how the framework relates to cognitive models of analogical reasoning, which have traditionally afforded a central role to the process of analogical mapping, driven by patterns of similarity over entities and relations [5, 69, 70, 71]. An intriguing possibility is that the relational bottleneck encourages neural networks to learn to implement such algorithms, by re-representing their inputs in terms of patterns of similarity (represented as inner products), though future work should aim to more precisely establish this link. Future work should also consider how the proposed framework relates to other theoretical perspectives on out-of-distribution generalization in neural networks, including group theoretic approaches [120, 121], as well as other cognitive processes relevant to abstraction, including attentional processes [122] and semantic cognition [123]. Additionally, much work has suggested that human reasoning is not purely relational, but instead depends on a mixture of concrete and abstract influences [124, 125, 126, 127]. This suggests the potential value of a more graded formulation that controls the amount of non-relational information allowed to pass through the bottleneck. Finally, the human capacity for abstraction surely depends not only on architectural biases such as those that we have discussed here, but also on the rich educational and cultural fabric that allows us to build on the abstractions developed by others [128]. In future work, it will be important to explore the interaction between education, culture and relational inductive biases.
@@
  • Can a more graded version of the relational bottleneck capture ‘content effects’ – in which abstract reasoning processes are influenced by the specific content under consideration, and therefore are not purely abstract or relational – while preserving a capacity for relational abstraction?
@@
  • What is the relationship between the relational bottleneck and traditional cognitive models of analogical reasoning? Does the relational bottleneck provide a useful inductive bias toward learning to implement processes such as analogical mapping?
  • How can other cognitive processes (attention, memory, etc.) be integrated with the relational bottleneck?
@@
  • How do architectural biases toward relational processing interact with the influence of training curricula and other cultural sources of abstraction (e.g., formal education)?
@@
In the context of deep learning, an approach that combines neural networks with separate external stores of information, typically with learnable mechanisms for writing to and reading from these stores, and in which retrieval is usually similarity-based (i.e., ‘content-addressable’). Often used to implement a form of episodic memory [132].
@@
Acknowledgements
AA is supported by funds provided by the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence). JDC is supported by Vannevar Bush Faculty Fellowship N00014-22-1-2002 from the Office of the Under Secretary of Defense for Research & Engineering, supported by ONR.